# On the generalization of language models from in-context learning and finetuning: a controlled study

### v2

@@ Discussion
In this paper, we performed controlled experiments on how language models generalize various types of novel information from in-context learning and finetuning. Overall, we find that the models generalize better on average along several dimensions from in-context learning. Using in-context learning to augment the finetuning dataset can exploit the complementary benefits of both to yield better performance.
@@
The distinct inductive biases of in-context learning and finetuning: A variety of works have studied the inductive biases of in-context learning. One common theme has been to emphasize that in-context learning can approximate gradient descent (e.g. Von Oswald et al., 2023), in settings where that behavior is optimal. However, a variety of other works have found that the inductive biases of in-context learning can vary depending on factors such as dataset diversity (Raventós et al., 2024) or model scale (Wei et al., 2023). Indeed, Elmoznino et al. (2024) argue that in-context learning Several works have explicitly noted the distinct inductive biases of in-context vs. in-weights learning (e.g. “rule-based” vs “exemplar-based” generalization) (Chan et al., 2022b; Shen et al., 2023; Awadalla et al., 2022)—while others have argued that they are more similar than they might seem (Mosbach et al., 2023). Our work contributes to this line of findings, but goes beyond the typical input-output tasks considered in these works, to consider other types of knowledge.
@@
Accessible information and learning by thinking: Lombrozo (2024) highlights how “learning by thinking” is a unifying theme across cognitive science and recent advances in AI—a system can acquire new information and skills purely through computation, without further inputs. Lombrozo highlights that while superficially this may seem paradoxical—information cannot be created—this further computation can increase the accessibility of information and thus improve performance in practice. This argument parallels the theoretical one made by Xu et al. (2020) on how computation can increase the accessibility of information. Our use of in-context inferences to improve finetuning performance beyond the original data follows this pattern. For example, the information about reversals and syllogisms is always hidden within the data, but finetuning on the in-context inferences make this information more explicit and thus more readily accessible at test time.
@@
Train-time inference scaling: Recently, various works have begun to explore test-time inference scaling to improve performance (e.g. Jaech et al., 2024; Guo et al., 2025). These findings complement prior studies exploring how scaling training compute (e.g. via larger models or more data) improves performance (e.g. Kaplan et al., 2020; Hoffmann et al., 2022). Our results illustrate that scaling train-time compute through in-context inference methods can help to improve some aspects of model generalization.
@@
Limitations: Our work suffers from several limitations. First, our main experiments rely on nonsense words and implausible operations. Although these counterfactual tasks allow us to avoid the possibility of dataset contamination, they may interfere with the performance of the models to some extent. For example, preliminary experiments (Appx. B.2) suggest that the ICL performance of the models on the Reversal Curse dataset degrades if the names are replaced with nonsense. Thus, tasks with more plausible entities may see greater benefits from ICL. It’s possible that finetuning makes the nonsense terms more familiar, and that this contributes to the gap between augmented finetuning and ICL. However, in that case we likely underestimate the gap between ICL and base finetuning (as the latter would effectively benefit from increased “familiarity” of the nonsense entities). Second, we have not experimented with other language models, which would enhance the generality of our results. However, since the individual phenomena we build upon — e.g. the reversal curse when finetuning (Berglund et al., 2024) vs. the ability to do reversals in context (e.g. Lampinen et al., 2024b) — have been documented across multiple models, we believe it is reasonable to extrapolate our results to other settings. However, future work should study more carefully the differences in how models learn and generalize in these settings, including newer reasoning models (e.g. Guo et al., 2025).
@@
Conclusions: We have explored the generalization of in-context learning and finetuning when LMs are exposed to various types of novel information structures. We find notable differences—often with ICL generalizing better for certain types of inferences—and propose ways of achieving better performance by adding in-context inferences to finetuning data. We hope this work will contribute to the science of understanding learning and generalization in foundation models, and the practicalities of adapting them to downstream tasks.
@@ B.2 Non-sensification and Long-context ICL
In the main paper, we observed that full-dataset in-context evaluation performs very well for the reversal task of Berglund et al. (2024) (cf. Fig. 2). However, for the nonsense reversals, the ICL performance is relatively low (cf. Fig. 3). We conjecture that this difference in performance can be attributed to the nonsense nouns used in the simple reversal task as an LLM does not have a good semantic prior over these nonsense nouns so it can’t utilize its long-context as effectively to do ICL. To test this hypothesis, we modify the dataset of Berglund et al. (2024) by nonsensifying all the celebrity names.
@@
Figure 6: Comparison of ICL with and without nonsensification of celebrity names in the Reversal Curse Paper (Berglund et al., 2024) dataset.
@@ B.3 Effect of Model Size
B.4 Efficacy of different prompts
@@ B.5 Process benchmark
This benchmark is built to test a model’s ability to learn a novel “process” from examples. Here, a process is a transformation of an input. This contrasts with learning facts or semantics (which is tested by the semantic structure benchmark above), given that processes may be learned and represented differently than facts (e.g., Geva et al., 2021; Ruis et al., 2024).
We target a few desiderata: (a) The processes are not familiar to the pretrained LM, requiring the model to go beyond mapping a familiar process to novel symbols, given that models can simply learn a latent mapping to that existing process (Treutlein et al., 2024). (b) At the same time, the task uses mostly familiar words and symbols, to avoid tokenization issues.
[…]
@@
Table 2: Derivatoid transformation rules for combinations of “primitive” expressions, where $u$ and $v$ are primitives, and $du$ and $dv$ are their derivatoids (as defined in Table 1)
@@ Methods
  • To understand whether the effect occurs at train-time or test-time, we also tried flipping the ordering of the responses so that the reflections came after the answers. E.g. “Answer: (sin(x)) * (35**x) + (x**35) * (meep(x)). Reflection: This looks like a Exponential function multiplied by a Trigonometric function.” This did not lead to the same improved performance. This may be because the augmentation effect is mostly a test-time effect (c.f. “chain of thought reasoning” (Wei et al., 2022)), or because the augmentations cannot affect representations in the right way at train-time, especially given the particular form of the augmentations “This looks like…”, which aren’t as sensical when they follow the final answer rather than the input.

### v3

@@ Discussion
In this paper, we performed controlled experiments on how language models generalize various types of novel information from in-context learning and finetuning. Overall, we find that the models generalize better on average along several dimensions from in-context learning. Using in-context reasoning to augment the finetuning dataset can exploit the complementary benefits of both to yield better performance.
@@
The distinct inductive biases of in-context learning and finetuning: A variety of works have studied the inductive biases of in-context learning. One common theme has been to emphasize that in-context learning can approximate gradient descent (e.g. vonoswald2023transformers), in settings where that behavior is optimal. However, a variety of other works have found that the inductive biases of in-context learning can vary depending on factors such as dataset diversity (raventos2024pretraining) or model scale (wei2023larger). Indeed, elmoznino2024context argue that in-context learning has an “Occam’s razor”-like bias towards minimizing complexity, which enhances its generalization. Several works have explicitly noted the distinct inductive biases of in-context vs. in-weights learning (e.g. “rule-based” vs “exemplar-based” generalization) (chan2022transformers; shen2023pretrained; awadalla2022exploring)—while others have argued that they are more similar than they might seem (mosbach-etal-2023-shot). Our work contributes to this line of findings, by moving beyond the typical input-output tasks considered in these works, to consider other types of knowledge—and illustrates some potential utility of these differences, by showing how the more flexible generalization of in-context learning can be exploited to improve the generalization of fine-tuning.
@@
Accessible information and learning by thinking: lombrozo2024learning highlights how “learning by thinking” is a unifying theme across cognitive science and recent advances in AI—a system can acquire new information and skills purely through computation, without further inputs. Lombrozo highlights that while superficially this may seem paradoxical—information cannot be created—this further computation can increase the accessibility of information and thus improve performance in practice. This argument parallels the theoretical one made by xu2020theory on how computation can increase the accessibility of information. Our use of in-context reasoning to improve finetuning performance beyond the original data follows this pattern. For example, the information about reversals and syllogisms is always hidden within the data, but finetuning on the in-context inferences make this information more explicit and thus more readily accessible at test time.
@@
Train-time inference scaling: Recently, various works have begun to explore test-time inference scaling to improve performance (e.g. jaech2024openai; guo2025deepseek). These findings complement prior studies exploring how scaling training compute (e.g. via larger models or more data) improves performance (e.g. kaplan2020scaling; hoffmann2022training). Our results illustrate that scaling train-time compute through in-context inference methods can likewise help to improve some aspects of model generalization.
@@
Limitations: Our work suffers from several limitations. First, our main experiments rely on nonsense words and implausible operations. Although these counterfactual tasks allow us to avoid the possibility of dataset contamination, they may interfere with the performance of the models to some extent. For example, preliminary experiments (Appx. B.2) suggest that the ICL performance of the models on the Reversal Curse dataset degrades if the names are replaced with nonsense. Thus, tasks with more plausible entities may see greater benefits from ICL. It’s possible that finetuning makes the nonsense terms more familiar, and that this contributes to the gap between augmented finetuning and ICL. However, in that case we likely underestimate the gap between ICL and base finetuning (as the latter would effectively benefit from increased “familiarity” of the nonsense entities). Second, we have not experimented with other language models, which would enhance the generality of our results. However, since the individual phenomena we build upon — e.g. the reversal curse when finetuning (berglund2024reversal) vs. the ability to do reversals in context (e.g. lampinen2024language) — have been documented across multiple models, we believe it is reasonable to extrapolate our results to other settings. However, future work should study more carefully the differences in how models learn and generalize in these settings, including newer reasoning models (e.g. guo2025deepseek).
Conclusions
@@
We have explored the generalization of in-context learning and finetuning when LMs are exposed to various types of novel information structures. We find notable differences—often with ICL generalizing better for certain types of inferences—and propose ways of achieving better performance by adding in-context reasoning traces to augment finetuning data. We hope this work will contribute to the science of understanding learning, reasoning, and generalization in foundation models, and the practicalities of adapting them to downstream tasks.
@@ A.1 Prompts
A.2 Datasets
Semantic structure
The full tree structure used to derive the benchmark has 110 entities structured into a tree with depth of up to 7, with branching factors ranging from 1-8, and numbers of node-specific properties ranging from 1-6.
The (non-nonsense) relations used, e.g. “are a type of” were converted into natural language by a range of templates (per relation). In turn, these were embedded in a larger range of phrasing and formatting structures designed to mimic some superficial details of an article about a topic.
An example training document is shown below.
Jund
Jund have yact. ("Phropt are a component of jund.") "Swug is a capability of jund." Jund are one kind of hipt. Jund can sunt; however, jund have the component splect.
* Thrund are the parent category of jund.
* (Jund are a subset of thont.)
* Jund have clud.
* Jund can glon.
* (Ust are a component of jund.)
* Spluff is a capability of jund.
Splect are a part of jund. Jund are a subset of thrund, and jund can swug.
Assuming this document were the only information in the training corpus, one example reversal QA pair that could be generated relative to it would be “Q: Clud are a part of…? A: Jund.”
A.3 Compute resources
Finetuning and evaluation experiments used TPUv4 and v5 resources. For the largest (semantic) dataset each finetuning pass took about 10-20 minutes on 128 devices; evaluating a finetuned model took about 10-20 minutes on 8 devices. ICL evaluation took about 1.5-2 hours on 64 devices, and augmentation took about 3 hours on 256 devices. As the other dataset are smaller, we estimate reproducing all experiments in the main text of this paper would take around 2000 TPU-hours.
@@ B.2 Non-sensification and Long-context ICL
In the main paper, we observed that full-dataset in-context evaluation performs very well for the reversal task of berglund2024reversal (cf. Fig. 2). However, for the nonsense reversals, the ICL performance is relatively low (cf. Fig. 3). We conjecture that this difference in performance can be attributed to the nonsense nouns used in the simple reversal task as an LLM does not have a good semantic prior over these nonsense nouns so it can’t utilize its long-context as effectively to do ICL. To test this hypothesis, we modify the dataset of berglund2024reversal by nonsensifying all the celebrity names.
@@
Figure 6: Comparison of ICL with and without nonsensification of celebrity names in the Reversal Curse Paper (berglund2024reversal) dataset.
@@ B.3 Effect of Model Size
B.4 Hyper-parameter Sweeps
We ablate batch size and training steps for the finetuning baseline. The results are reported in Fig. 8 and Fig. 9. The numbers reported in the main paper are based on the strongest performing baseline.
Figure 8: Batch size ablation on the semantic structure benchmark on the Finetuning baseline. The performance of batch size of 16 is slightly better so we use that in the main paper. However, changing the batch size doesn’t improve the performance of the finetuning baseline significantly and the gains we see for the ICL and Augmentation methods in the main paper are not due to the poor choice of batch size selection for the baseline.
Figure 9: The effect of varying the training steps in the Finetuning baseline on the semantic structure benchmark. Overall the results are fairly stable across the training steps. Note that changing the number of training steps doesn’t improve the performance of the finetuning baseline significantly and the gains we see for ICL and Augmentation methods in the main paper are not due to the under or over-training of the baseline.
B.5 Efficacy of different prompts
@@ B.6 Process benchmark
This benchmark is built to test a model’s ability to learn a novel “process” from examples. Here, a process is a transformation of an input. This contrasts with learning facts or semantics (which is tested by the semantic structure benchmark above), given that processes may be learned and represented differently than facts (e.g., geva_transformer_2021; ruis_procedural_2024).
We target a few desiderata: (a) The processes are not familiar to the pretrained LM, requiring the model to go beyond mapping a familiar process to novel symbols, given that models can simply learn a latent mapping to that existing process (treutlein2024connecting). (b) At the same time, the task uses mostly familiar words and symbols, to avoid tokenization issues.
To fulfill these desiderata, we designed “derivatoids”, which are a derivative-like transformation of math expression. “Primitive” expressions are transformed according to Table 1, and combinations of primitives are transformed according to to Table 2. For example, the input (log(x,39))(64x)(log(x,39))*(64**x) should be converted to (64x)(log(x,39))+(log39log39(x))(x64)(64**x)*(log(x,39))+(log39log39(x))*(x**64).
@@
Table 2: Derivatoid transformation rules for combinations of “primitive” expressions, where $u$ and $v$ are primitives, and dudu and dvdv are their derivatoids (as defined in Table 1)
@@ Methods
  • To understand whether the effect occurs at train-time or test-time, we also tried flipping the ordering of the responses so that the reflections came after the answers. E.g. “Answer: (sin(x)) * (35**x) + (x**35) * (meep(x)). Reflection: This looks like a Exponential function multiplied by a Trigonometric function.” This did not lead to the same improved performance. This may be because the augmentation effect is mostly a test-time effect (c.f. “chain of thought reasoning” (wei_chain_2022)), or because the augmentations cannot affect representations in the right way at train-time, especially given the particular form of the augmentations “This looks like…”, which aren’t as sensical when they follow the final answer rather than the input.